### v1

# Down with the Hierarchy: The ‘H’ in HNSW Stands for “Hubs”

Abstract.
Driven by recent breakthrough advances in neural representation learning, approximate near-neighbor (ANN) search over vector embeddings has emerged as a critical computational workload. With the introduction of the seminal Hierarchical Navigable Small World (HNSW) algorithm, graph-based indexes have established themseves as the overwhelmingly dominant paradigm for efficient and scalable ANN search. As the name suggests, HNSW searches a layered hierarchical graph to quickly identify neighborhoods of similar points to a given query vector. But is this hierarchy even necessary? A rigorous experimental analysis to answer this question would provide valuable insights into the nature of algorithm design for ANN search and motivate directions for future work in this increasingly crucial domain. To that end, we conduct an extensive benchmarking study covering more large-scale datasets than prior investigations of this question. We ultimately find that a flat graph retains all of the benefits of HNSW on high-dimensional datasets, with latency and recall performance essentially identical to the original algorithm but with less memory overhead. Furthermore, we go a step further and study why the hierarchy of HNSW provides no benefit in high dimensions, hypothesizing that navigable small world graphs contain a well-connected, frequently traversed “highway” of hub nodes that maintain the same purported function as the hierarchical layers. We present compelling empirical evidence that the Hub Highway Hypothesis holds for real datasets and investigate the mechanisms by which the highway forms. The implications of this hypothesis may also provide future research directions in developing enhancements to graph-based ANN search.
1. Introduction
Near neighbor search is a fundamental problem in computational geometry that lies at the heart of countless practical applications. From industrial-scale recommendation (Feng et al., 2022) to retrieval-augmented generation (Lewis et al., 2020) and even to computational biology (Zhao et al., 2024), numerous data-intensive tasks utilize similarity search at some location in the stack. As a result, similarity indexes are very well-studied (Guo et al., 2019; Malkov and Yashunin, 2016; Johnson et al., 2017; Aguerrebere et al., 2023; Subramanya et al., 2019) with multiple large-scale benchmarks and leaderboards to compare techniques (Aumüller et al., 2018; Simhadri et al., 2022).
Historically, the state-of-the art for near neighbor search involved constructing sophisticated tree-based data structures, such as $kd$-trees (Bentley, 1975) and cover trees (Beygelzimer et al., 2006), that guaranteed exact solutions while avoiding a brute-force examination of all points. However, the recent advent of large-scale neural representation learning, including large language models (LLMs), placed a significant strain on these classical methods that were developed to target a much lower-dimensional search space. In response, the community has turned to approximate search methods. While alternative approximate indexing methods such as locality-sensitive hashing (Indyk and Motwani, 1998) and product quantization (Jegou et al., 2010), have garnered significant interest, graph-based approaches generally achieve the strongest performance on established ANN benchmarks (Aumüller et al., 2018; Simhadri et al., 2022). Introduced in 2016, the Hierarchical Navigable Small World (HNSW) algorithm (Malkov and Yashunin, 2016), emerged as one of the first high-performance graph-based search indexes at scale and still enjoys immense popularity to this day with over 4300 Github stars [^1] https://github.com/nmslib/hnswlib.
As the name implies, a core feature of the HNSW index is its hierarchically layered graph akin to a skip list (Pugh, 1990) where a search iteratively traverses through graphs of increasing density before converging to a neighborhood of similar points in the final graph layer. By drawing intuition from skip lists, the HNSW authors argue that the initial coarse graph layers allow for efficiently identifying the neighborhood of similar points in the collection through fewer overall comparisons.
Despite the immense popularity of HNSW, however, the algorithm suffers from multiple scalability bottlenecks. For instance, the hierarchical layers introduce a hefty memory overhead. Moreover, Malkov and Yashunin (2016) also note that the hierarchy can reduce the overall system throughput in distributed settings when compared to a flat graph.
While the high cost of this hierarchy in HNSW is traditionally justified by the low latency of graph-based indexes, we ask the question of whether the hierarchy is even necessary in the first place. A growing body of results suggest that the hierarchy may be a vestigial artifact left behind by the lower-dimensional problems of the past. For example, Lin and Zhao (2019) published a set of experiments suggesting that the hierarchical component of HNSW is only beneficial for low dimensional inputs ($d<32$). Other recent works have observed similar problems with the hierarchy. For example, Coleman et al. (2022) present an ablation study showing the same behavior in the appendix of their paper. This behavior seems to be an increasingly well-known “folklore” of the similarity search community, but there is currently an absence of a thorough investigation and analysis of this phenomenon.
An exhaustive benchmark comparison on the utility of the HNSW hierarchy would go a long way towards improving our understanding of graph-based similarity search in high-dimensional space and inform future research directions in developing improvements. However, there are several key difficulties with executing such a benchmarking study. The first issue is that running a meaningful ablation study requires first engineering a similarity search implementation over a flat graph that reaches performance parity with HNSW and other state of the art approaches. This is an increasingly nontrivial task, as the community has invested significant performance engineering efforts into these codebases such that a non-expert implementation has little hope of being competitive.
The second challenge is designing benchmarking experiments in a manner that avoids the confounding effect of the performance of particular software implementation when making conclusions about an algorithm’s efficacy. Finally, benchmarking near-neighbor search can be highly nuanced. For example, it is possible that hierarchical structures are strongly beneficial for isolated slices of the input. These may sharply reduce the 99th percentile latency while not affecting the median.
Perhaps most importantly, we still have no satisfactory understanding of why hierarchy does not help. Hierarchical structures are a mainstay of algorithm design, where a common trick is to reduce an $O(n)$ search process to a sublinear one by traversing a (balanced) hierarchy (Pugh, 1990; Guibas and Sedgewick, 1978; Mikolov et al., 2013; Cormen et al., 2022). Arguably, it is counterintuitive for this idea to fail to hold in the context of high-dimensional similarity search – especially when we have strong positive results that hierarchy helps in low dimensions (Beygelzimer et al., 2006; Dolatshah et al., 2015; Ram and Sinha, 2019; Lin and Zhao, 2019). Thus, an exhaustive benchmark and deeper analysis into the necessity of the hierarchy in HNSW would shed further light on the nature of algorithm design in high-dimensional spaces and thus may be of independent interest to the community as well.
Figure 1. We hypothesize that in high dimensions, graph-based $k$-NN indices naturally form a “highway-feeder” structure, where a small subset of nodes and edges are easily reached, well-connected, and heavily traversed.
1.1. Contributions
In this paper, we take on the question of whether the hierarchical component of HNSW is truly necessary, or if we can achieve the same performance on large-scale benchmarks with simply a flat navigable small world graph. In particular, we make the following contributions:
Benchmarking the hierarchy: We rigorously benchmark HNSW to understand whether the hierarchy is necessary. To do so, we reproduce and extend the hierarchy ablations of previous studies, finding that, on high-dimensional vector datasets, it is indeed beneficial to remove the ‘H’ from HNSW.
Why does hierarchy not help? We hypothesize that the hierarchy benefits decrease due to hubness. Hubness is a high-dimensional phenomenon that causes a skewed distribution in the near-neighbor lists of search queries. We hypothesize that hubness leads to preferential attachment in the similarity search graph, inducing the formation of easily-traversed highways that connect disparate regions of the graph. This hypothesis, which we call the hub-highway hypothesis, explains why we no longer need the hierarchy in high dimensions; we can simply traverse the intrinsic highway structure that naturally forms in high-dimensional spaces. To that end, we conduct a series of experiments to investigate the hub-highway hypothesis. Our results ultimately support this hypothesis, showing that hubness is responsible for driving the connectivity of similarity search graphs. This insight opens up exciting new research directions in graph construction, link pruning, and graph traversal. Our work also reveals that the hub-highway is an intrinsic and naturally-forming structure in high-dimensional proximity graphs, which we believe has important implications for the scaling potential of many traversal heuristics. [^2] Specifically, that the benefits of sophisticated initialization methods will decay under hub-style preferential attachment.
Our specific contributions are as follows.
  • We release an implementation for a flattened version of HNSW [^3] https://github.com/nmslib/hnswlib, called FlatNav [^4] https://github.com/BlaiseMuhirwa/flatnav, that reaches performance parity with the original version with considerable memory savings.
  • We demonstrate that hierarchy does not improve performance in either the median or tail latency case by building HNSW and FlatNav indexes over 13 popular benchmark datasets ranging in size from 1 million to 100 million vectors.
  • We conduct an analysis of hubness phenomena in high-dimensional metric spaces and the resulting HNSW graphs, finding strong empirical support for the hub-highway hypothesis.
Takeaway: Our benchmarks reveal that HNSW can be significantly optimized for modern high-dimensional embedding workloads. In particular, our implementation saves roughly $38\%$ of peak memory consumption compared to hnswlib (and sizable further headroom is likely). Our results confirm the folklore of the similarity search community, conclusively demonstrating that we can remove the hierarchy on high-dimensional inputs with impunity.
2. Background: Similarity Search & HNSW
In the similarity search (or $k$-NNS) problem, we are interested in retrieving $k$ elements from a dataset $\mathcal{D}=\{\mathbb{x}_{i},\ldots,\mathbb{x}_{n}\}\subset\mathbb{R}^{d}$ that minimize the distance to a given query $q\in\mathbb{R}^{d}$ (or, equivalently, maximize the vector similarity). More precisely, given a similarity function $\phi:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R},$ the nearest neighbor $\mathbb{x}^{*}\in\mathcal{X}$ of $q$ is defined as
$\mathbb{x}^{*}\coloneqq\operatorname*{arg\,max}_{\mathbb{x}_{i}\in\mathcal{D}}\phi(\mathbb{x}_{i},q)$
where $\phi$ is usually the $\ell_{2}$ or cosine distance. With the enormity of modern data workloads and the underlying vector dimensionality, it becomes computationally infeasible to exhaustively search for the true top-$k$ neighbors for any query $q.$ Thus, approximate search algorithms trade-off quality of the search for lower latency.
In the approximate nearest neighbor search (ANNS) regime, we evaluate the quality of the search procedure typically by the Recall@$k$ metric. More formally, suppose a given ANNS search algorithm outputs a subset $\mathcal{O}\subseteq\mathcal{D},|\mathcal{O}|=k,$ and let $G\subseteq\mathcal{D}$ be the true $k$ nearest neighbors of a query $q.$ We define Recall@$k$ by $\frac{|\mathcal{O}\cap G|}{k}$. ANNS algorithms seek to maximize this metric while retrieving results as quickly as possible.
2.1. HNSW Overview
With this formalization of the ANNS problem, we will now briefly review the key elements of the HNSW algorithm, which is the central focus of our benchmarking study. As we alluded to previously, HNSW builds off of prior work in navigable small world graph indexes which was introduced in (Malkov et al., 2014). Small world graphs are a well-studied phenomenon in both computing and the social sciences and are primarily defined by the fact that the average length of a shortest path between two vertices is small (typically scaling logarithmically with the number of nodes in the network) (Travers and Milgram, 1977; Watts and Strogatz, 1998; Kleinberg, 2000). Small world graphs are also often characterized by the presence of well-connected hub nodes which we discuss further in the next section.
While small world graphs are, by construction, suited for efficient greedy graph traversal, the HNSW authors argue that the polylogarithmic scaling of the search process is still too inefficient for the demands of near neighbor search on large datasets. This claim motivates the design of HNSW where the hierarchy allows for computing a fixed number of distances in each graph layer independent of the network size.
Specifically, the HNSW index is constructed in an iterative fashion. For a newly inserted element $x$, the algorithm will randomly select a maximum layer $l$ and then insert the new point into every layer up to $l$. This randomized process is executed with an exponentially decaying probability distribution such that, in expectation, each subsequent layer has exponentially more nodes than its predecessor. Within a layer, HNSW greedily adds edges between $x$ and its $ef$ closest neighbors (where $ef$ is a hyperparameter) where the neighbors consist of previously inserted points. This process then repeats in the subsequent layer below using the closest neighbors found in the prior graph as entry points. Through this process, the top layer of the hierarchy will be the coarsest directed graph, consisting of the fewest nodes and edges, and the bottom layer will be the densest and contain all of the nodes, each with connections to $ef$ number of neighbors. As an additional, and important, optimization, HNSW also implements the pruning heuristic of (Arya and Mount, 1993) that will prune an edge from $u$ to $v$ if there exists another edge from $u$ to a neighbor $w$ of $v$ such that the distance from $u$ to $w$ is less than that of $u$ to $v$.
The search procedure of HNSW also executes iteratively where the algorithm maintains a list of candidate points at each layer of the hierarchical graph before returning the final list $k$ nearest neighbor candidates after traversing the base graph layer.
3. Related Work
3.1. Near-Neighbor Benchmarks
Due to the importance of the $k$-NNS problem, there have been several large-scale near-neighbor benchmarks in recent years. The original ANN Benchmarks  (Aumüller et al., 2018) made significant impact because it provided the first standard evaluation of near-neighbor search algorithms. Since its inception, the benchmark has grown in scope from a handful of algorithms to include over 30 methods on 9 datasets. However, most of the datasets included in ANN Benchmarks are relatively small by modern standards at around one million points. To reflect the growing need for high-dimensional, large-scale embedding search, NeurIPS hosted a competition track known as Big ANN Benchmarks (Simhadri et al., 2022), which evaluates algorithms on five billion-scale datasets. Graph-based indexes easily and consistently place in among the top three algorithms in each benchmark. However, these benchmarks are not sufficient for our purposes because they focus exclusively on overall throughput metrics and do not report tail latency numbers such as the 99th percentile, which is important in our case for teasing apart the impact of hierarchy in HNSW.
Hierarchy Studies: Hierarchical structures are considered a fundamental component of graph-based near-neighbor algorithms. As a result, numerous high-performing algorithms use some form of hierarchy including HNSW (Malkov and Yashunin, 2016), ONNG (Iwasaki and Miyazaki, 2018) PANNG (Iwasaki, 2016), and HCNNG (Munoz et al., 2019). However, this practice has recently come under scrutiny. Dobson et al. (2023) show that the hierarchy may be unnecessary for some workloads, based on the observation that HNSW under-performs both HCNNG (which uses a shallower hierarchy) and DiskANN (Subramanya et al., 2019) (which does not use a hierarchy). Lin and Zhao (2019) present results on synthetic data showing that the hierarchy is only beneficial for low-dimensional data ($d<32$), but analyze only a limited number of small real-world datasets and – most importantly – do not explain why the hierarchy fails to provide value. In this work, we aim to address these gaps in the literature through more comprehensive benchmarks coupled with an analysis of why hierarchical structures in high-dimensional space may not add value. In the following section, we aim to reproduce the results of both Malkov and Yashunin (2016) and Lin and Zhao (2019) to confirm that we can independently replicate their findings with our own software implementation before proceeding to new experiments on larger-scale benchmark datasets.
3.2. Hubness in High Dimensional Spaces
Astute readers might observe that the HNSW graph construction algorithm described above does not explicitly enforce the small world property and instead adds edges between nodes based on their proximity in the metric space. The connection between proximity and the small world property arises due to hubness.
Specifically, hubness is a property of high-dimensional metric spaces where a small subset of points (the “hubs”) occur a disproportionate number of times in the near-neighbor lists of other points in the dataset (Radovanovic et al., 2010). In other words, a small fraction of nodes are highly connected to other points.
A related, but slightly different, notion of hubness that has been at the center of theoretical analysis is concentration of distances in high-dimensional spaces. It has been shown that, in expectation, the $\ell_{2}$ distance between independent and identically distributed (i.i.d) vectors grows with $\sqrt{d}$ while the variance tends to a constant as $d$ approaches infinity (Talagrand, 1994). As a result, the $\ell_{2}$ distance loses its discriminative power as $d$ increases. In fact, this concentration phenomenon is not only constrained to $\ell_{2}$ but also applies to other $\ell_{p}$ and even fractional norms (François et al., 2007).
Due to undesirable consequences of the hubness phenomenon, such as poor clustering quality, a large body of work has focused on hubness reduction strategies. For instance, (Zelnik-Manor and Perona, 2004) introduced local scaling which scales distances $d(\mathbb{x},\mathbb{y})$ by accounting for local neighborhood information. Interestingly, our work stands in contrast to this literature on hubness reduction by presenting a case study where hubs provide tangible value in an algorithmic setting, namely in accelerating greedy traversal in near neighbor proximity graphs. This result may be of independent interest to machine learning and algorithms researchers as well.
4. Reproduction of Prior Studies
In this section, we present two replicability studies of prior experimental evaluations on the efficacy of HNSW versus a non-hierarchical NSW index. As we discussed in the previous section, these prior works possess limitations in experimental design, scope of benchmarking datasets, and a lack of analysis into understanding the results, which motivates our work in this paper. Nevertheless, we use these prior studies as a starting point to see if we can independently replicate these results via our own FlatNav implementation. Such a reproduction would both further validate the soundness of these previous experiments over the test of time as well as provide validation of the correctness of FlatNav before we proceed to new, larger-scale benchmarks.
4.1. Reproduction of Malkov and Yashunin (2016)
For our first replicability study, we look to reproduce one of the core experimental result of the original HNSW paper (Malkov and Yashunin, 2016). To validate that the hierarchical layers provide a speedup over a flat NSW index, Malkov and Yashunin benchmark their HNSW implementation (available through the hnswlib library) against the sw-graph NSW implementation. Furthermore, we also add our own flat graph implementation via our flatnav library as an additional comparison. Following the original HNSW paper, we generate two synthetic random hypercube datasets consisting of 4 and 8 dimensions, respectively. In particular, each dimension of these random vectors is sampled uniformly at random from $[0,1)$. We report the results of this replicability study in Figure 2 where we find that HNSW indeed achieves superior latency than flatnav and sw-graph on these low-dimensional datasets.
We also had the opportunity to discuss our findings with the lead author of (Malkov and Yashunin, 2016) who confirmed that the hierarchy provides a robust speedup on these low-dimensional datasets but noted the performance on higher dimensional vectors remained less clear, which further motivated us to take up the benchmarking study in the next section.
Figure 2. Mean Latency vs. Recall of HNSW, SW-Graph and FlatNav across dimensions $d=4,8$. We find that HNSW achieves lower mean latency than its flat index counterparts in FlatNav and SW-Graph, matching the findings of Malkov and Yashunin (2016)
4.2. Reproduction of Lin and Zhao (2019)
We now present our results reproducing the findings of Lin and Zhao (2019) in their 2019 benchmarking study, which concluded that the hierarchical structure of HSNW is only beneficial for low-dimensional data. Although Lin and Zhao primarily focus on synthetic, randomly generated datasets, we looked to see if we could independently replicate these findings before turning our attention to benchmarking on modern, large-scale embedding datasets in the next section.
Following the same setup as (Lin and Zhao, 2019), we generate a series of random vector datasets of varying dimensionality where each vector component is sampled uniformly at random from the range $[0,1)$. In particular, we consider dimensionalities of $d=4,8,16$ and 32. As in (Lin and Zhao, 2019), we set the number of near neighbors to retrieve to $k=1$ (departing from the default of $k=100$ we use elsewhere in this paper).
Figure 3. Median Latency vs. Recall of HNSW and FlatNav across dimensions $d=4,8,16,32$. We observe that the hierarchical structure accelerates search only when $d<32$, matching the findings of Lin and Zhao (2019)
In summary, we successfully replicated the experiments benchmarking HNSW versus a flat graph from two prior research papers. Notably, both of these previous works primarily experiment with randomly generated vector data with very low dimensionality by the standards of modern machine learning. Coupled with our findings in the next section where we find no discernible difference between HSNW and a flat graph index on high-dimensional datasets, our results suggest a simple decision criterion for selecting a search index: For dimensionality $d<32$, HNSW and the hierarchy provide a speedup. Otherwise, the simplicity and memory savings of a flat index provide more benefit.
5. FlatNav Benchmarking Experiments
In this section, we report the results of our benchmarking study comparing the performance of flat HNSW search to hierarchical search on a suite of standard high-dimensional benchmark datasets drawn from real machine learning models.
A challenge in designing such benchmarking experiments is separating the confounding impact of performance engineering in the implementation of an approach from any purely algorithmic advantages of a method. We argue that this is a pitfall that has clouded several past benchmarking studies in ANN search. To circumvent this challenge, we fix the implementation in our experimental design such that the same code is used in constructing the base navigable small world graph. In particular, we use the popular hnswlib library as our baseline HNSW implementation. To benchmark the flat index performance, we extract the bottom graph layer from hnswlib and reimplement the HNSW’s greedy search heuristic over this flat graph. To the best of our knowledge, we provide the first benchmarking study on the utility of HNSW’s hierarchy that makes an effort to achieve parity in the implementations of the two techniques to center the focus on the central algorithmic question.
5.1. Software Performance Optimizations
One major challenge in prior near neighbor benchmarking studies is the confounding effect of software performance optimizations when making conclusions about algorithmic efficacy. To address this issue, our implementation of a flat similarity search graph in flatnav closely follows the performance optimizations native to hnswlib. We use SIMD vector instructions to efficiently compute the distance between any two vectors, which is the key operation underpinning similarity search. Depending on the hardware type, we provide distance computations using SSE, AVX and AVX-512 intrinsics for float, uint8_t and int8_t data types. The choice of which intrinsics family to use is a compile-time choice, which is determined based on the dataset dimensionality and hardware support, similar to hnswlib. For all benchmark experiments, we use the float data type.
Furthermore, one common problem in constructing and querying graph-based indexes is the random data access pattern which can hamper latency and throughput. To improve search throughput, hnswlib implements prefetching using the _mm_prefetch intrinsic which reduces memory access latency by prefetching data into the L1 cache. Since caching can significantly affect latency metrics, we control for this performance optimization by leveraging the same intrinsic in the flatnav implementation so that our benchmarks strictly adhere to algorithmic differences.
5.2. Datasets and Compute
We utilize the benchmark datasets released through the popular leaderboards ANN Benchmarks (Aumüller et al., 2018) and Big ANN Benchmarks (Simhadri et al., 2022). The specific datasets and their associated statistics are presented in Table 1. For the Big ANN Benchmark datasets, we consider both the 10M and 100M collection of vectors for which the ground truth near neighbors have previously been computed and released. We did not experiment with the largest Big ANN datasets with 1 billion vectors since constructing HNSW indexes at this scale requires over 1.5TB of RAM, which exceeded our compute resources. Below, we include our benchmarking results for the four 100M-scale datasets available through Big ANN Benchmarks. We see that our flat HNSW implementation achieves performance parity with the hierarchical HNSW implementation.
Dataset Dimensionality Collection Size Num Queries
BigANN 128 100M 10K
Microsoft SpaceV 100 100M 29.3K
Yandex DEEP 96 100M 10K
Yandex Text-to-Image 200 100M 100K
GloVe {25, 50, 100, 200} 1.2M 10K
NYTimes 256 290K 10K
GIST 960 1M 1K
SIFT 128 1M 10K
MNIST 784 60K 10K
DEEP1B 96 10M 10K
Table 1. Dataset Statistics. The datasets marked by $\dagger$ are from the BigANN benchmarks (Simhadri et al., 2022). The remaining are taken from ANN Benchmarks (Aumüller et al., 2018)
For our benchmarks on datasets consisting of fewer than 100M vectors in the collection, we use an AWS c6i.8xlarge instance with an Intel Ice Lake processor and 64GB of RAM. We selected this particular public cloud instance to facilitate accessible reproducibility of our experiments. For the 100M-sized large-scale experiments, we use a cloud server equipped with an AMD EPYC 9J14 96-Core Processor and 1 TB of RAM.
5.3. Latency Results
5.3.1. BigANN Benchmarks (Simhadri et al., 2022)
In Figures 4 and 5, we compare latency metrics for HNSW and FlatNav at the 50th and 99th percentile for the four 100M datasets from BigANN benchmarks listed in Table 1. All of our results support the conclusion that flatnav achieves nearly identical performance to hnswlib.
From the results in Figures 4 and 5, we observe that there is no consistent and discernable gap between FlatNav and HNSW in both the median and tail latency cases. These results suggest that the hierarchical structure of HNSW provides no tangible benefit on practical high-dimensional embedding datasets.
Figure 4. p50 Latency vs. Recall for HNSW and FlatNav over four 100M-scale benchmark datasets with dimensionality between 96 and 200. We observe that the flat index achieves essentially identical performance to the hierarchical HNSW index.
Figure 5. p99 Latency vs. Recall for HNSW and FlatNav over four 100M-scale benchmark datasets with dimensionality between 96 and 200. As with the median case, we observe nearly identical performance between the two indexes.
5.3.2. ANN Benchmarks (Aumüller et al., 2018)
Moreover, we repeat the same experimental setup comparing HNSW and FlatNav on the ANN Benchmark datasets listed in Table 1. In Figures 6 and 7, we report the p50 and p99 latency of all of the non-GloVe ANN Benchmarks. Although these datasets are smaller in scale than the BigANN Benchmarks, we still see no discernible difference in latency between HNSW and FlatNav which supports our hypothesis that the vector dimensionality and not the size of the collection is the main driver of eliminating the need for hierarchical search in small world graphs. We see further evidence of this idea in Table 8, where we plot the median latency versus recall on the four GloVe benchmark datasets ranging in dimensionality from 25 to 200. Ultimately, these results, coupled with our reproducibility studies in the previous section, provide compelling evidence that for high-dimensional datasets, which are the standard in modern similarity search workloads, there is no apparent performance benefit to the hierarchical layers of HNSW.
Figure 6. p50 Latency vs. Recall for HNSW and FlatNav over five ANN Benchmark datasets ranging in dimensionality from 96 to 960. The flat index again achieves practically indistinguishable performance from the hierarchical graph index.
Figure 7. p99 Latency vs. Recall for HNSW and FlatNav over five ANN Benchmark datasets ranging in dimensionality from 96 to 960.
Figure 8. p50 Latency vs. Recall for HNSW and FlatNav over the GloVe ANN Benchmark datasets with respective dimensionalities 25, 50, 100, and 200. Again, we see no clear difference between the two indexes in terms of performance.
5.4. Memory
In this section we measure the memory savings from removing the hierarchy by running a memory profiler during index construction. In terms of memory allocation, similarly to flatnav, hnswlib allocates static memory during index construction comprising base layer node allocation, a visited node list and a list of mutexes in a multi-threaded setting. Additionally, it also incurs memory cost attributable to the hierarchy, particularly maintaning the dynamically allocated links between nodes at each layer.
We benchmarked both libraries against the BigANN 100M dataset shown in Table 1 and traced memory consumption during the index construction stage parallelized over $32$ cores. We observe that, at its peak, hnswlib reaches $183$ GBs of RAM while flatnav peaks at $113$ GBs, demonstrating up to $38\%$ reduction in peak memory consumption. This shows that we are able to save significant memory by removing the hierarchy, and it is likely that we can optimize flatnav implementation to save memory further.
One caveat of these reported memory savings is that we are comparing different two software implementations in hnswlib and flatnav. Since hnswlib is a mature library widely used by practitioners as well as researchers, it supports more features than flatnav and thus must maintain additional complexity whereas our implementation, while performant, is more of a research prototype. Therefore, differences in code may account for a significant part of the peak memory usage differences. Nevertheless, we believe our findings are relevant and still noteworthy given that hnswlib is so widely adopted. By demonstrating that we can considerably reduce the memory overhead of hnswlib without sacrificing performance, we hope to bring the community’s attention to the opportunities for further optimization in this direction.
6. The Hub Highway Hypothesis
In the previous section, we established through a series of careful benchmarking experiments that there appears to be no discernible difference in performance between HNSW and its flat counterpart with no hierarchical graph structure. We now turn our attention to studying why the hierarchy appears to provide no benefit in the search process.
In our experiments, we observed that a small fraction of nodes appear in the set of near neighbors for a disproportionate number of other vectors. We thus conjecture that the hub structure prevalent in high-dimensional data performs the same functional role as the hierarchy.
Hypothesis 0 (Hub Highways).
In high-dimensional metric spaces, $k$-NN proximity graphs form a highway routing structure where a small subset of nodes are well-connected and heavily traversed, particularly in the early stages of graph search.
We remark that the existence of hub nodes in high-dimensional space is not a new observation and has in fact been studied in the prior literature on small world graphs. The novelty of our hypothesis lies in connecting the idea of hubness to the notion of accelerating near neighbor search in proximity graphs. In particular, we conjecture that near neighbor queries over proximity graphs in high dimensions often spend the majority of their time visiting hub nodes early on in the search process before converging to a local neighborhood of near neighbors. This procedure succeeds because hub nodes are very well connected to other parts of the graph and thereby efficiently route queries to the appropriate neighborhood in much the same manner that the layered hierarchy purports to do.
In many applications, such as clustering, hubness is often considered an undesirable property that harms algorithmic performance, thereby motivating the need for hubness reduction techniques (Zelnik-Manor and Perona, 2004). On the contrary, our Hub Highway Hypothesis presents a case where hubness offers a distinct advantage in a practical algorithmic setting.
In the remainder of this section, we present an experimental design and a series of results that provide empirical evidence in the affirmative for the existence of such a highway routing mechanism amongst hubs in navigable small world graphs.
6.1. Methodology
In this section, we will present a formal definition and characterization of hubness and present empirical results in support of our hypothesis.
Following (Radovanovic et al., 2010), let $\mathbb{x},\mathbb{x}_{1},\ldots,\mathbb{x}_{n}$ be a collection of vectors drawn from the same probability distribution with support $\mathcal{S}\subseteq\mathbb{R}^{d}$, and let $\phi:\mathcal{S}\times\mathcal{S}\to\mathbb{R}$ be a distance function.
Furthermore, for $1\leq i,k\leq n$, let $p_{i,k}$ be defined by
$p_{i,k}(\mathbb{x})=\begin{cases}1&\text{if $\mathbb{x}$ is among the $k$-NN set of $\mathbb{x}_{i}$ under $\phi$}\\ 0&\text{otherwise }\end{cases}$
Now let $N(\mathbb{x})$ be the random variable defined by
$N(\mathbb{x})\coloneqq\sum_{i=1}^{n}p_{i,k}(\mathbb{x})$
which represents the number of vectors that have $\mathbb{x}$ included in their $k$-nearest neighbors.
For any dataset $\mathcal{D}$, we can compute $N(\mathbb{x}),\mathbb{x}\in\mathcal{D}$, which yields a discrete distribution $N_{k}$. We are interested in the skewness of this distribution, given by
$S_{N_{k}}=\frac{\mathbb{E}\left[\left(N_{k}-\mu_{N_{k}}\right)^{3}\right]}{\sigma^{3}_{N_{k}}}$
This measure characterizes the asymmetry of the $k$-occurrence distribution $N_{k}$, and it is the metric most often used to estimate the presence of hubs. Thus, in our experiments we use $S_{N_{k}}$ to quantify the presence of hubness in a particular dataset. The more skewed the distribution of $N_{k}$, the more dataset $\mathcal{D}$ tends to have clusters of vectors (hubs) that keep regularly occurring in the $k$-NN sets of other vectors.
Table 2. Datasets
Dataset Dimensionality Collection Size Num Queries
GIST 960 1M 1k
GloVe 100 1.2M 10k
NYTimes 256 290K 10k
Yandex-DEEP 96 10M 10k
Microsoft-SpaceV 100 10M 29.3k
IID Normal {16, 32, 64, 128, 256, 1024, 1536} 1M 10k
IID Normal {16, 32, 64, 128, 256, 1024, 1536} 1M 10k
Synthetic and ANN Benchmark Datasets: To illustrate evidence for the hub-highway hypothesis, we experiment with both real and synthetic datasets as shown in Table 2. In addition to a subset of ANN datasets, we generate synthetic datasets by drawing vectors from the standard normal distribution. The choice of the normal distribution is important because, as (Low et al., 2013) observe, data sampled from the normal distribution exhibit a stronger hubness phenomenon due to strong density gradient than other distributions, such as sampling uniformly from a $d$-dimensional hypercube. Furthermore, since the hubness phenomenon is accentuated in high-dimensional spaces, we generate random data with increasing dimensionality which allows us to study the effect of these intrinsic highway structures that form as $d$ increases.
In the following three experiments we show empirical evidence supporting the hub-highway hypothesis. We start by examining the distribution of the number of times each node in the similarity search index is visited during search given a fixed number of queries, which we denote as $P_{m}(\mathbb{x}_{i})$ to indicate that node $\mathbb{x}_{i}$ is visited exactly $m$ times during a fixed number of $k$-NN queries for each one of the datasets in Table 2. We show that this distribution is skewed to the right, thus confirming that certain nodes (highway nodes) are visited a disproportionate number of times. Next, using this distribution, we selectively choose the most visited nodes in the similarity search graph to be the hub node clusters and show evidence that these nodes are more connected to each other than random nodes using hypothesis tests. In the last experiment, we show that not only are these highway nodes more connected, they also allow for faster graph traversals for queries, hence supporting the claim that we no longer need hierarchy in high dimensional vector search where fast query times evolve as a result of the presence of the hub-highways.
6.2. Skewness of the Node Access Distribution
As described above, for this study we consider the discrete distribution $P_{m}(\mathbb{x}_{i})$, $\mathbb{x}_{i}\in\mathcal{D}$, where $\mathcal{D}$ is one of the datasets in Table 2. For all these experiments we construct a HNSW index with parameters listed in Table 3 and import the base layer of the graph into FlatNav to construct a flat index.
Table 3. Similarity search index parameters
$m$ $ef$-construction $ef$-search $k$
32 100 200 100
(a) Cosine KDE Plot
(b) L2 KDE Plot
Figure 9. Log-normalized node access count distribution $P_{m}(\mathbb{x}_{i})$ plotted with kernel smoothing (Kernel Density Estimation). (a) Distribution for different datasets that use cosine similarity along with the skewness $\tilde{\mu}_{3}$ of $P_{m}(\mathbb{x}_{i})$. (b) Distribution for the datasets that use $l_{2}$ distance and the corresponding skweness values.
Figure 9 shows the log-normalized node access count distribution for different datasets. We observe that as the dimension $d$ increases, for $\ell_{2}$ distance-based datasets this distribution becomes more skewed to the right, indicating that a subset of nodes are visited more often than the rest of the nodes in the graph. It is worth noticing that even for $d\in\{1024,1536\}$, the cosine distance function exhibits anti-hub properties, effectively forcing little to no hubness in the data. This is consistent with the results observed by (Radovanovic et al., 2010). This increased skewness of the $P_{m}(\mathbb{x}_{i})$ distribution as $d$ increases explains how highway nodes occur as a result of the intrinsic dimensionality of the datasets.
6.3. Subgraph Connectivity of the Hub-Highway Nodes
Having shown how highway nodes occur, in the next set of experiments we show empirical evidence confirming that these nodes exhibit stronger connectivity in the graph than any given random nodes in the similarity search graph. We set up a hypothesis test as follows with the null hypothesis that hub nodes are not any more connected to each other than randomly chosen nodes. Let $\mathbb{x}_{i},1\leq i\leq n$, be the vectors in a similarity search index for any dataset $\mathcal{D}$ under consideration.
[…]
  • For each identified hub node in the index, we compute the number of hub nodes adjacent to it. This yields a discrete distribution of the number of hub nodes to which each hub node is connected.
  • Similarly, we select a set of random nodes in the set $V\coloneqq\mathcal{D}\setminus\bigcup_{i=1}^{n}h(\mathbb{x}_{i})=1$. For each node $\mathbb{x}_{i}\in V$, we compute the number of hub nodes adjacent to it to get an equivalent distribution.
We set up both a parametric test with a two-sample t-test and a non-parametric test using the Mann-Whitney U-test (Mann and Whitney, 1947), which is suitable particularly for the ANN datasets since it makes no normality assumption. A well-known consequence of the Central Limit Theorem (CLT) (Durrett, 2019) is that in non-parametric statistical tests like the Mann-Whitney U test, as the sample size increases, the U-statistic converges to a normal distribution. This, in turn, leads to $p$-values that approach 0 whenever there is a true difference between any two populations under consideration. To avoid this pitfall of the $p$-value, we also consider effect size in our test and constrain the sample size to be 1000 for both populations.
Table 4 shows the $p$-value results in the regime where we use the p95 of the node access distribution to select hub nodes. Assuming the regular 0.05 threshold for rejecting the null hypothesis, notice that in all cases but 5 we reject the null hypothesis under the Mann-Whitney $U$-test (and all cases but 6 for the two-sample $t$-test). Unsurprisingly, for the synthetic datasets, the effect size is much higher for $l$-2 datasets since this distance metric induces stronger hubness in the dataset, effectively increasing the likelihood that the subgraph induced by hub nodes is more connected. It is worth noting that the Yandex-DEEP benchmark does not exhibit a similar hub node subgraph connectivity effect seen in the majority of the other benchmarks. Yandex-DEEP embeddings were generated from the last fully-connected layer of GoogleNet (Szegedy et al., 2014) by applying PCA and $l$-2 normalization. In this case $l$-2 normalization alleviates the effects of hubs since the magnitude of the vectors does not disproportionately affect distance computations.
Similarly, Table 5 shows the $p$-value results in the case we use the p99 threshold of the node access counts to select hub nodes. With the 0.05 statistical significance threshold we are able to rejec the null hypothesis in all cases but 1 for both the Mann-Whitney $U$-test and the two-sample $t$-test. In this case we realize much higher effect sizes, which is not surprising since the p99 threshold selects an even more refined set of hub nodes. Thus, the two statistical tests show strong evidence that hub nodes form more connected subgraphs than those induced by randomly sampling nodes in the similarity search graph.
Table 4. Two-sample T-test and Mann-Whitney U-test p-values. Hub nodes are selected using the p95 threshold for the node access distribution.
Dataset Dim Mann-Whitney Two-Sample T-Test Effect Size IID Normal (Angular) 16 0.3629 0.3090 0.0267 IID Normal (L2) 16 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.3737 IID Normal (Angular) 32 0.0335 0.0516 0.0872 IID Normal (L2) 32 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.4275 IID Normal (Angular) 64 0.0216 0.0148 0.1165 IID Normal (L2) 64 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.3965 IID Normal (Angular) 128 0.0083 0.0083 0.1284 IID Normal (L2) 128 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.3773 IID Normal (Angular) 256 0.0009 0.0007 0.1723 IID Normal (L2) 256 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.2620 IID Normal (Angular) 1024 0.1000 0.1114 0.0652 IID Normal (L2) 1024 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.2361 IID Normal (Angular) 1536 0.0957 0.1141 0.0645 IID Normal (L2) 1536 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.2512 GloVe 100 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.2550 NYTimes 256 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.4488 GIST 960 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.3645 Yandex-DEEP 96 0.5002 0.5000 0.0000 Microsoft-SpaceV 100 0.1586 0.1585 0.0535
Table 5. Two-sample T-test and Mann-Whitney U-test p-values. Hub nodes are selected using the p99 threshold for the node access distribution.
Dataset Dim Mann-Whitney Two-Sample T-Test Effect Size
IID Normal (Angular) 16 0.0006 0.0006 0.1745
IID Normal (L2) 16 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.6621
IID Normal (Angular) 32 0.0347 0.0347 0.0972
IID Normal (L2) 32 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.8173
IID Normal (Angular) 64 0.0359 0.0417 0.0927
IID Normal (L2) 64 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.8725
IID Normal (Angular) 128 0.0093 0.0070 0.1316
IID Normal (L2) 128 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.8428
IID Normal (Angular) 256 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.3110
IID Normal (L2) 256 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.8582
IID Normal (Angular) 1024 0.1472 0.1318 0.0598
IID Normal (L2) 1024 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.8314
IID Normal (Angular) 1536 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.2356
IID Normal (L2) 1536 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.8568
GloVe 100 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.7642
NYTimes 256 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.9305
GIST 960 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.6829
Yandex-DEEP 96 0.0013 0.0013 0.1614
Microsoft-SpaceV 100 0.0011 0.0011 0.1644
6.4. Hub-Highway Nodes Allow Queries to Traverse the Graph Faster
In this section we show that not only are highway nodes more connected to each other than random nodes, they also enable fast traversals along the similarity search graph. Intuitively this is not surprising given our results that these nodes are more connected to each other and have long range connections between them that allow queries to navigate across distant regions in navigable small world graphs efficiently. To measure this effect, we construct similarity search indexes and for each query $q$, track the sequence of nodes visited before beam search terminates.
Suppose $\mathbb{x}_{1},\mathbb{x}_{2},\ldots,\mathbb{x}_{l}$ is a length-$l$ sequence of such nodes for query $q.$ We use the hub node assignment heuristic discussed in section 6.3 to get a label $h(\mathbb{x}_{i})$ for each node in the sequence. We then split the sequence into bins and compute the percentage of hub nodes visited in each bin. For bin $B_{i}$, this is given by
$\left(\frac{1}{|B_{i}|}\right)\sum_{\mathbb{x}_{j}\in B_{i}}h(\mathbb{x}_{j})$
where $|B_{i}|$ is the bin size. We average this value over all queries and fix the bin size to $30$. Figure 10 shows results for Gist, GloVe, Microsoft SpaceV and the Yandex Deep benchmark datasets. We observe that earlier in the search procedure, queries tend to concentrate in the highway structures as shown by the percentage of hub nodes visited in the earlier bins. Since highway nodes are able to connect disparate regions in the graph, this result confirms that they propel queries to navigate the similarity search graph faster as they are visited the most earlier in the search. Worth noting is that for GloVe, as shown in Figure 10(b), the percentage of hub nodes visited is much lower across all the bins. This is consistent with prior findings that show that the cosine similarity exhibit anti-hub properties as it is not affected by the magnitude of the vectors compared with the $l_{2}$ distance.
(a) Gist
(b) GloVe
(c) Microsoft SpaceV
(d) Yandex Deep
Figure 10. Highway nodes allow queries to traverse the graph faster.
6.5. Discussion
We observe substantial evidence from the above sequence of experiments supporting the Hub Highway Hypothesis. Although it has long been established that the hubness phenomenon negatively affect common applications, such as clustering and even near neighbor search recall, existing ANNS methods, such as HNSW and NSG (Zhao et al., 2023) mostly emphasize algorithmic improvements and performance optimizations only. We have shown that leveraging the inherent structures in the data, particularly hub-highway occurrences, should be central to the design of new scalable similarity search indexes.
As discussed earlier, the skewness of the $k$-occurrence distribution $N_{k}$ directly affects the quality of search in terms of recall. Given our empirical results showing that hub-highways allow queries to traverse the graph faster, interesting future research directions might consider leveraging this distribution to improve edge pruning heuristics or judiciously choose which edges to explore during beam search over proximity graphs.
7. Conclusion
Approximate near neighbor search has become an increasingly crucial computational workload in recent years with the seminal Hierarchical Navigable Small World (HNSW) algorithm continuing to garner significant interest and adoption from practitioners. Given the impact of HNSW on both the research community and production systems in industry, a deep understanding of the algorithm’s inner workings would be of immediate value.
To that end, this paper presents the first comprehensive study on the utility of the hierarchical component of HNSW over numerous large-scale datasets and performance metrics. Ultimately, we find that the hierarchy of HNSW provides no clear benefit on high-dimensional datasets and can be removed without any discernible loss in performance while providing memory savings.
While similar observations have been made before in the literature (Lin and Zhao, 2019; Coleman et al., 2022; Dobson et al., 2023), we are, to our knowledge, the first to conduct an exhaustive study over modern benchmark datasets and taking extensive care to compare implementations with performance engineering parity. Furthermore, we go beyond prior works and study why the hierarchy does not help, culminating in our introduction of the Hub Highway Hypothesis, an empirical result on how proximity graphs built over high-dimensional metric spaces leverage a small subset of well-connected nodes to traverse the network quickly, akin to highways and exits in transportation systems.
As part of our benchmarking study and analysis, we also release a high-performance flat navigable small world implementation, called FlatNav, that we hope will serve as an additional resource to both researchers and industry practitioners. We also hope that our work can inspire future directions in algorithm design for ANN search, such as in better strategies for exploiting the hub highway structure we observed or even designing new algorithms that do manage to leverage the idea of hierarchy effectively. In the short term, we believe that our results provide immediate implications for practitioners seeking to save memory or simplify their vector database implementations and we look forward to further partnering with the community on these endeavors.
Acknowledgements.
We thank Yury Malkov, the lead author of (Malkov and Yashunin, 2016) and maintainer of the hnswlib library, for valuable insights into the original HNSW experimental evaluation as well as thorough feedback on this manuscript.